# Modular transformations of on-shell actions of (root-)$\overline{\text{T}}$ deformed holographic CFTs

### v2

@@ Appendix B BTZ black holes with complex periods
It is straightforward to transform the metric (4.1) into the FG gauge (2.11) with
@@
From the holographic stress tensor (A.22), one can derive the holographic mass and momentum
@@
The entropy of the black hole is equal to
@@
which implies that the periods are transformed as
@@
where $\omega_{1}=2\pi,\omega_{2}=0,\omega_{1}+\text{i}\omega_{2}=\bm{\omega},$ and $\beta_{1}=\theta,\beta_{2}=\beta,\beta_{1}+\text{i}\beta_{2}=\bm{\beta}$. Noting the this transformation is a combination of Weyl transformation and rotation, the holographic mass and momentum should transform accordingly as
@@
which lead to (4.5) and (4.6). In addition, these two quantities correspond to the conserved charges of $\partial_{\hat{\phi}}$ and $\partial_{\hat{t}}$ respectively.
@@ Appendix C On-Shell Action of TMG Theory
In this appendix, we will derive the on-shell action (5.6) of TMG theory by assuming the Smarr relation still holds. The Smarr relation can be viewed as a relationship between energy and the entropy and it holds separately for the holomorphic and anti-holomorphic sectors. Therefore, it is not modified even when $c_{L}\neq c_{R}$. Starting from the following expressions
@@
here we have chosen a real $\omega$. Actually, with the help of (4.5) and (4.6), it is straightforward to obtain the result with complex period $\bm{\omega}$ as follows
@@
Next by applying $L_{0}=-\frac{\pi^{2}}{\bm{\omega}^{2}}$ and $\bm{\beta}=\bm{\omega}\bm{\tau}$, the above result can be simplified to

### v3

@@ Appendix B BTZ black holes with complex periods
In this appendix, we find the mass, angular momentum and entropy of a rotating BTZ black hole in terms of the complex periods of the torus on which the dual CFT lives. To do so, we first transform the BTZ metric (4.1) into the FG gauge (2.11). It is straightforward to verify that
@@
Then, from the holographic stress tensor (A.22), one can derive the holographic mass and momentum
@@
Moreover, the entropy of the black hole is equal to
@@
which implies that the periods are transformed as follows
@@
where $\omega_{1}=2\pi,\omega_{2}=0,\omega_{1}+\text{i}\omega_{2}=\bm{\omega},$ and $\beta_{1}=\theta,\beta_{2}=\beta,\beta_{1}+\text{i}\beta_{2}=\bm{\beta}$. Noting that this transformation is a combination of Weyl transformation and rotation, the holographic mass and momentum should transform accordingly as
@@
which leads to (4.5) and (4.6). In addition, these two quantities correspond to the conserved charges of $\partial_{\hat{\phi}}$ and $\partial_{\hat{t}}$, respectively.
@@ Appendix C On-Shell action of TMG theory
In this appendix, we will derive the on-shell action (5.6) of TMG theory by assuming the Smarr relation still holds. The Smarr relation can be viewed as a relationship between the energy and entropy and it holds separately for the holomorphic and anti-holomorphic sectors. Therefore, it is not modified even when $c_{L}\neq c_{R}$. Starting from the following expressions
@@
Here we have chosen a real $\omega$. Actually, by applying (4.5) and (4.6), it is straightforward to obtain the result with complex period $\bm{\omega}$ as follows
@@
Next, by applying $L_{0}=-\frac{\pi^{2}}{\bm{\omega}^{2}}$ and $\bm{\beta}=\bm{\omega}\bm{\tau}$, the above result can be simplified to